### v2

# Almost All Vectorial Functions Have Trivial Extended-Affine Stabilizers

Abstract
We prove that asymptotically almost all vectorial functions over finite fields have trivial extended-affine stabilizers. As a consequence, the number of EA-equivalence classes is asymptotically equal to the naive estimate, namely the total number of functions divided by the size of the EA-group, with vanishing relative error. Furthermore, we derive upper bounds on collision probabilities for both extended-affine and CCZ equivalences. For EA-equivalence, we leverage the trivial-stabilizer result to establish a matching lower bound, yielding a tight asymptotic formula that shows two independently sampled functions are EA-equivalent with super-exponentially small probability. The results validate random sampling strategies for cryptographic primitive design and show that functions with nontrivial EA-stabilizers form an exponentially rare subset.
Keywords: Boolean functions, extended-affine equivalence, EA-stabilizer, group action, orbit counting, asymptotic enumeration
1 Introduction
Vectorial Boolean functions play a crucial role in symmetric cryptography, serving as S-boxes in block ciphers and hash functions. The classification of these functions is typically studied modulo certain equivalence relations that preserve important cryptographic properties. Extended-affine (EA) equivalence is one of the most fundamental such relations. It preserves differential uniformity, a key measure of resistance against differential cryptanalysis, as well as the Walsh spectrum and algebraic degree. A closely related notion is CCZ-equivalence, introduced by Carlet, Charpin, and Zinoviev [4], which also preserves some cryptographic properties and is more general than EA-equivalence.
A substantial body of recent work has focused on the algorithmic and structural aspects of EA- and CCZ-equivalence. Canteaut and Perrin [3] showed that CCZ-equivalence can be decomposed into EA-equivalence and an operation called twisting, and derived upper bounds on the number of EA-equivalence classes within a single CCZ-equivalence class. Budaghyan, Calderini, and Villa [1] introduced sufficient conditions for determining when CCZ-equivalence is strictly more general than EA-equivalence combined with taking inverses of permutations. On the algorithmic side, Kaleyski [9] surveyed invariants for EA- and CCZ-equivalence of APN and AB functions and evaluated their effectiveness in distinguishing inequivalent instances. Canteaut, Couvreur, and Perrin [2] presented an efficient algorithm for recovering the EA-equivalence transformation between quadratic functions, along with a new invariant based on the ortho-derivative. These works address the problem of testing whether two given functions are EA-equivalent.
Two fundamental questions arise in the study of EA-equivalence:
  1. 1.
    Classification: How many EA-equivalence classes are there? Since EA-equivalent functions share the same cryptographic properties, counting equivalence classes determines the number of essentially distinct functions relevant for symmetric cryptography. This is crucial for understanding the diversity of cryptographic primitives.
  2. 2.
    Random sampling: When searching for functions with desirable properties via random generation, is there a risk of repeatedly sampling EA-equivalent functions? In other words, does random search remain effective modulo EA-equivalence, or do collision events significantly limit the exploration of the equivalence class space?
Both questions are intimately connected to the structure of EA-stabilizer groups. The EA-stabilizer of a function $F$ is the set of all pairs of affine permutations $(A_{\rm in},A_{\rm out})$ satisfying $A_{\rm out}\circ F\circ A_{\rm in}=F$; it measures how much “self-equivalence” the function possesses. By the orbit-stabilizer theorem, the size of each equivalence class is inversely proportional to the stabilizer size, which directly impacts both the total number of classes and the collision probability. However, explicit computation of stabilizers is infeasible except for very small dimensions, as the function space grows doubly exponentially with the input dimension.
1.1 Our Contributions
This paper resolves these questions by establishing that asymptotically almost all vectorial functions have trivial EA-stabilizers. Our main contributions are as follows:
  1. 1.
    Main theorem (Theorem 3.4): We prove that for a uniformly random vectorial function over $\mathbb{F}_{q}^{n}\to\mathbb{F}_{q}^{m}$, the probability of having a nontrivial EA-stabilizer decays super-exponentially.
  2. 2.
    Counting equivalence classes (Corollary 3.7): We establish that the number of EA-equivalence classes is asymptotically equal to the total number of functions divided by the size of the EA-group, with vanishing relative error. This provides a simple and conceptually clear asymptotic formula grounded in the triviality of stabilizers.
  3. 3.
    Collision probabilities (Proposition 3.8 and Corollary 3.9): We derive upper bounds on collision probabilities for both EA- and CCZ-equivalence, and show that for EA-equivalence, the upper bound is asymptotically tight. This establishes that two independently sampled functions are EA-equivalent with super-exponentially small probability, validating random sampling strategies for cryptographic primitive design.
Collectively, these results demonstrate that functions with nontrivial EA-stabilizers form an exponentially rare subset. This complements existing approaches focused on specific cryptographic properties and provides a theoretical foundation for understanding the prevalence of nontrivial stabilizers in the space of vectorial functions.
1.2 Related Work
Previous work on stabilizers of Boolean functions has primarily focused on simpler group actions. Clausen [6] proved that almost all Boolean functions (not vectorial) have trivial stabilizers under the action of $\mathrm{GL}(n,2)$. Clote and Kranakis [7] established similar results for invariance groups under variable permutations. However, these results apply only to Boolean functions $\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ and do not address EA-equivalence, which involves affine transformations on both domain and range, a key distinction for vectorial functions.
A different line of research focuses on counting equivalence classes. Hou [8] obtained exact formulas for affine equivalence classes and asymptotic formulas for EA-equivalence classes of Boolean functions using Burnside’s lemma and compound matrix rank estimates. Lu et al. [10] extended these results to vectorial Boolean functions, deriving exact formulas via Möbius inversion and asymptotic formulas by applying Hou’s matrix-theoretic techniques.
Our work unifies these two lines of research by establishing the connection between stabilizers and orbit enumeration for vectorial functions under EA-equivalence. We prove that almost all vectorial functions have trivial EA-stabilizers, extending the trivial-stabilizer results of Clausen and Clote-Kranakis to the vectorial EA setting. As an immediate consequence via the orbit-stabilizer theorem, we recover the asymptotic orbit counts of Hou [8] and Lu et al. [10] with an elementary probabilistic proof, and reveal the underlying structural reason: the naive estimate (total number of functions divided by the group size) is asymptotically exact because stabilizers are trivial with high probability.
1.3 Paper Organization
Section 2 reviews necessary background on group actions, boolean functions and EA-equivalence. Section 3 presents our main results, including the probabilistic analysis of EA-stabilizers and the asymptotic count of EA-classes. Section 4 concludes with a discussion of implications and open problems.
2 Preliminaries
2.1 Group Actions and Probabilistic Tools
Let $q$ be a prime power and let $\mathbb{F}_{q}$ denote the finite field with $q$ elements. Let $U=\mathbb{F}_{q}^{n}$ and $W=\mathbb{F}_{q}^{m}$ denote finite vector spaces over $\mathbb{F}_{q}$. The general linear group $\mathrm{GL}(n,q)$ consists of all invertible $n\times n$ matrices over $\mathbb{F}_{q}$, and has order
$|\mathrm{GL}(n,q)|=\prod_{i=0}^{n-1}(q^{n}-q^{i}).$
Definition 2.1 (Affine general linear group).
The affine general linear group $\mathrm{AGL}(U)$ is the semidirect product $U\rtimes\mathrm{GL}(n,q)$, acting on $U$ by $\sigma(x)=Px+a$ with $P\in\mathrm{GL}(n,q)$ and $a\in U$. Its order is $|\mathrm{AGL}(U)|=q^{n}\cdot|\mathrm{GL}(n,q)|$.
Definition 2.2 (Orbits and stabilizers).
Let $G$ be a group acting on a set $X$. For $x\in X$, the $G$-orbit of $x$ is
$G\cdot x:=\{g\cdot x:g\in G\}.$
The stabilizer of $x$ is
$\mathrm{Stab}_{G}(x):=\{g\in G:g\cdot x=x\}.$
For each $g\in G$, the fixed-point set is
$\mathrm{Fix}(g):=\{x\in X:g\cdot x=x\}.$
Lemma 2.3 (Orbit-stabilizer theorem).
For any $x\in X$, the size of the orbit $G\cdot x$ satisfies
$|G\cdot x|=\frac{|G|}{|\mathrm{Stab}_{G}(x)|}.$
Lemma 2.4 (Burnside’s lemma).
The number of $G$-orbits is given by
$|X/G|=\frac{1}{|G|}\sum_{g\in G}|\mathrm{Fix}(g)|.$
We use standard asymptotic notation: for functions $f,g:\mathbb{N}\to\mathbb{R}_{>0}$, we write $f(n)=O(g(n))$ if there exist constants $C>0$ and $n_{0}$ such that $f(n)\leq C\cdot g(n)$ for all $n\geq n_{0}$. We write $f(n)=\Omega(g(n))$ if $g(n)=O(f(n))$, and $f(n)=o(g(n))$ if $\lim_{n\to\infty}f(n)/g(n)=0$.
Lemma 2.5 (Union bound).
If $E_{1},\dots,E_{k}$ are events, then
$\mathbb{P}\Bigl(\bigcup_{i=1}^{k}E_{i}\Bigr)\leq\sum_{i=1}^{k}\mathbb{P}(E_{i}).$
2.2 Vectorial Functions and Equivalence Relations
Definition 2.6 (Vectorial functions).
We write $\mathcal{F}=\{F:U\to W\}$ for the set of all functions from $U$ to $W$. Note that $|\mathcal{F}|=q^{mq^{n}}$.
When $q=2$, functions in $\mathcal{F}$ are called vectorial Boolean functions. They play a central role in symmetric cryptography, where they are used as S-boxes in block ciphers and hash functions. The cryptographic quality of a vectorial function is measured by several properties. Differential uniformity quantifies resistance to differential cryptanalysis and is defined as the maximum number of solutions to F(x+a)F(x)=bF(x+a)-F(x)=b over all nonzero $a$ and all $b$. The Walsh spectrum and nonlinearity measure resistance to linear cryptanalysis, while the algebraic degree controls the complexity of algebraic attacks. Functions optimizing these properties, such as almost perfect nonlinear (APN) and almost bent (AB) functions, are of particular interest. Both EA-equivalence and CCZ-equivalence (defined below) preserve these cryptographic properties. For further details on these properties and their cryptanalytic significance, we refer to [5].
Definition 2.7 (Extended-affine equivalence).
The extended-affine (EA) group
$\Gamma:=\mathrm{AGL}(U)\times\mathrm{AGL}(W)$
acts on $\mathcal{F}$ by conjugating the input and output with affine maps:
$(A_{\rm out},A_{\rm in})\cdot F:=A_{\rm out}\circ F\circ A_{\rm in}.$
Explicitly, if $A_{\rm in}(x)=Px+a$ and $A_{\rm out}(y)=Qy+b$, then
$\bigl((P,a,Q,b)\cdot F\bigr)(x)=Q\,F(Px+a)+b.$
Two functions $F,G\in\mathcal{F}$ are EA-equivalent, denoted by $F\simeq_{\mathrm{EA}}G$, if they lie in the same $\Gamma$-orbit. The EA-stabilizer of $F$ is $\mathrm{Stab}_{\Gamma}(F)$.
Definition 2.8 (Graph of a function).
For a function $F:U\to W$, the graph of $F$ is the subset
$\Gamma_{F}:=\{(x,F(x)):x\in U\}\subseteq U\times W=\mathbb{F}_{q}^{n+m}.$
Definition 2.9 (CCZ-equivalence).
Two functions $F,G:U\to W$ are CCZ-equivalent (Carlet-Charpin-Zinoviev equivalent) if their graphs $\Gamma_{F}$ and $\Gamma_{G}$ are affine equivalent as subsets of $U\times W$. That is, $F$ and $G$ are CCZ-equivalent, denoted by $F\simeq_{\mathrm{CCZ}}G$, if there exists $L\in\mathrm{AGL}(U\times W)$ such that $L(\Gamma_{F})=\Gamma_{G}$.
CCZ-equivalence is more general than EA-equivalence: every pair of EA-equivalent functions is also CCZ-equivalent.
3 Main Results
3.1 Fixed-Point and Orbit Structure of Affine Permutations
Before analysing EA-stabilizers, we establish key properties of affine permutations over $\mathbb{F}_{q}$ that will control the orbit structure.
Lemma 3.1 (Fixed-point structure).
Let $\sigma(x)=Px+a$ be an affine permutation of $U=\mathbb{F}_{q}^{n}$ with $(P,a)\neq(I,0)$. Then the fixed-point set
$\{x\in U:\sigma(x)=x\}$
is either empty or an affine subspace of dimension at most n1n-1. In particular, $|\{x\in U:\sigma(x)=x\}|\leq q^{n-1}$.
Proof.
The fixed-point equation $Px+a=x$ is equivalent to (PI)x=a(P-I)x=-a. Since this is a system of linear equations, the solution set (if nonempty) is an affine subspace. If $P\neq I$, then $\operatorname{rank}(P-I)\geq 1$, so the solution space has dimension at most n1n-1. If $P=I$ and $a\neq 0$ (pure translation), there are no fixed points. ∎
Lemma 3.2 (Orbit count).
Let $\sigma$ be any permutation of a finite set of size $N$, with $f$ fixed points and $s$ orbits. Then
$s\leq\frac{N+f}{2}.$
In particular, if $\sigma(x)=Px+a$ is a nontrivial affine permutation of $U=\mathbb{F}_{q}^{n}$, then the number of orbits satisfies
$s\leq\frac{q^{n}+q^{n-1}}{2}=\frac{q+1}{2q}\cdot q^{n}.$
Proof.
The $f$ fixed points each form singleton orbits. The remaining NfN-f non-fixed points lie in orbits of size at least $2$, so they contribute at most (Nf)/2(N-f)/2 orbits. Hence
$s\leq f+\frac{N-f}{2}=\frac{N+f}{2}.$
For a nontrivial affine permutation, Lemma 3.1 gives $f\leq q^{n-1}$, so
$s\leq\frac{q^{n}+q^{n-1}}{2}=\frac{q^{n-1}(q+1)}{2}=\frac{q+1}{2q}\cdot q^{n}.$
3.2 Probabilistic Analysis of EA-Stabilizers
We now analyse the stabilizer distribution and deduce precise asymptotics for the number of EA-classes.
Lemma 3.3 (Fixed-point bound for EA-group elements).
Let $g=(P,a,Q,b)\in\Gamma$ be nontrivial. Then the number of $F\in\mathcal{F}$ satisfying $g\cdot F=F$ is bounded by
$|\mathrm{Fix}(g)|\;\leq\;q^{cmq^{n}}$
for some constant $c<1$ independent of $n,m$.
Proof.
The fixed-point condition $g\cdot F=F$ reads
$Q\,F(Px+a)+b=F(x)\quad(\forall x\in U).$ (1)
We distinguish two cases.
Case 1: $(P,a)\neq(I,0)$. Let $\sigma(x):=Px+a$; by assumption $\sigma\neq\mathrm{id}$. Lemma 3.2 shows that $\sigma$ has at most
$s\leq\frac{q+1}{2q}\cdot q^{n}$
orbits. Decompose $U$ into $\sigma$-orbits:
$U=\bigsqcup_{i=1}^{s}O_{i},\qquad O_{i}=\{x_{i},\sigma(x_{i}),\sigma^{2}(x_{i}),\dots,\sigma^{L_{i}-1}(x_{i})\},$
where $L_{i}\geq 1$ is the length of orbit $O_{i}$.
Fix an orbit $O_{i}$ of length $L_{i}$. Iterating (1) starting from $x_{i}$ gives
$F(x_{i})$ $=Q\,F(\sigma(x_{i}))+b$
$=Q^{2}F(\sigma^{2}(x_{i}))+(I+Q)b$
$\;\;\vdots$
$=Q^{L_{i}}F(x_{i})+\sum_{j=0}^{L_{i}-1}Q^{j}b.$
Hence $F(x_{i})$ must satisfy
$\bigl(I-Q^{L_{i}}\bigr)F(x_{i})=\sum_{j=0}^{L_{i}-1}Q^{j}b.$ (2)
The number of solutions $F(x_{i})\in W$ to (2) is at most $q^{\dim\ker(I-Q^{L_{i}})}\leq q^{m}$. Once $F(x_{i})$ is chosen, the values $F(\sigma^{k}(x_{i}))$ for $1\leq k<L_{i}$ are uniquely determined by (1).
Repeating this argument for each orbit, we see that $F$ is determined by choosing at most one free parameter (of size $\leq q^{m}$) per orbit. Since $s\leq(q+1)/(2q)\cdot q^{n}$, the total number of admissible $F$ is at most
$|\mathrm{Fix}(g)|\leq(q^{m})^{s}\;\leq\;q^{m\cdot\frac{q+1}{2q}\cdot q^{n}}=q^{\frac{q+1}{2q}mq^{n}}.$
Hence $c=(q+1)/(2q)$ suffices in this case.
Case 2: $(P,a)=(I,0)$ but $(Q,b)\neq(I,0)$. Now (1) becomes
$Q\,F(x)+b=F(x)\quad\Longleftrightarrow\quad(I-Q)F(x)=b\qquad(\forall x\in U).$
Since $(Q,b)\neq(I,0)$, either $Q\neq I$ or $b\neq 0$. If $Q\neq I$, then $\operatorname{rank}(I-Q)\geq 1$, so $\dim\ker(I-Q)\leq m-1$; the equation (IQ)F(x)=b(I-Q)F(x)=b has at most $q^{m-1}$ solutions for each $x$. If $Q=I$ and $b\neq 0$, there are no solutions. In either case, for each $x$ independently, $F(x)$ has at most $q^{m-1}$ choices. Hence
$|\mathrm{Fix}(g)|\;\leq\;(q^{m-1})^{q^{n}}=q^{(m-1)q^{n}}.$
Thus c=11/mc=1-1/m works when $m\geq 1$.
Combining both cases, we may take $c=\max\bigl\{(q+1)/(2q),1-1/m\bigr\}<1$. ∎
Theorem 3.4 (Almost all functions have trivial stabilizers).
For a uniformly random $F\in\mathcal{F}$,
$\mathbb{P}\bigl(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\bigr)\;\leq\;|\Gamma|\cdot q^{-(1-c)mq^{n}}\;=\;q^{-\Omega(mq^{n})}.$
Proof.
Observe that $\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}$ if and only if there exists a nontrivial $g\in\Gamma$ such that $g\cdot F=F$. For each such $g$, define the event
$E_{g}:=\{F\in\mathcal{F}:g\cdot F=F\}=\mathrm{Fix}(g).$
By Lemma 3.3, we have
$\mathbb{P}(E_{g})=\frac{|\mathrm{Fix}(g)|}{|\mathcal{F}|}\leq q^{-(1-c)mq^{n}}.$
Applying the union bound over all $|\Gamma|-1$ nontrivial elements of $\Gamma$, we obtain
$\mathbb{P}\bigl(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\bigr)$ $=\mathbb{P}\Bigl(\bigcup_{g\in\Gamma\setminus\{\mathrm{id}\}}E_{g}\Bigr)$
$\leq\sum_{g\in\Gamma\setminus\{\mathrm{id}\}}\mathbb{P}(E_{g})$
$\leq(|\Gamma|-1)\cdot q^{-(1-c)mq^{n}}$
$\leq|\Gamma|\cdot q^{-(1-c)mq^{n}}.$
To complete the proof, we bound $|\Gamma|$. Recall that
$|\mathrm{GL}(n,q)|=\prod_{i=0}^{n-1}(q^{n}-q^{i})\leq(q^{n})^{n}=q^{n^{2}},$
so $|\mathrm{AGL}(U)|=q^{n}\cdot|\mathrm{GL}(n,q)|\leq q^{n^{2}+n}=q^{O(n^{2})}$. Similarly, $|\mathrm{AGL}(W)|=q^{O(m^{2})}$. Hence
$|\Gamma|=|\mathrm{AGL}(U)|\cdot|\mathrm{AGL}(W)|\leq q^{O(n^{2}+m^{2})},$
and the result follows. ∎
In the binary case ($q=2$), Lemma 3.2 gives $(q+1)/(2q)=3/4$, so for $m\geq 4$ we have $c=\max\{3/4,1-1/m\}=3/4$. The probability bound becomes
$\mathbb{P}\bigl(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\bigr)\;\leq\;2^{O(n^{2}+m^{2})}\cdot 2^{-(1/4)m2^{n}}\;=\;2^{-\Omega(m2^{n})}.$
This probability decays super-exponentially in $n$, demonstrating that vectorial Boolean functions with nontrivial EA-stabilizers are exponentially rare. In other words, cryptographic S-boxes generically possess trivial EA-stabilizers: the overwhelming majority of functions in $\mathcal{F}$ have trivial stabilizer groups, and only an exponentially negligible fraction admit any nontrivial element in their EA-stabilizer.
Example 3.5 (Concrete bounds for cryptographic parameters).
We evaluate the bound from Theorem 3.4 for parameters commonly used in symmetric cryptography ($q=2$, $c=3/4$).
  1. 1.
    44-bit S-boxes ($n=m=4$): The group size satisfies $\log_{2}|\Gamma|=\log_{2}|\mathrm{AGL}(4,2)|^{2}=2(4+\log_{2}|\mathrm{GL}(4,2)|)=2(4+\log_{2}20160)<2\cdot 19=38$, and the exponent in the decay is $(1-c)m2^{n}=\frac{1}{4}\cdot 4\cdot 16=16$. Hence
    $\mathbb{P}\bigl(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\bigr)\;\leq\;2^{38-16}=2^{22},$
    which does not yield a nontrivial bound. This is expected: the function space $|\mathcal{F}|=2^{64}$ is not large enough relative to the group size for asymptotic freeness to manifest.
  2. 2.
    88-bit S-boxes ($n=m=8$): We have $\log_{2}|\Gamma|\leq 2(8^{2}+8)=144$, and the exponent in the decay is $(1-c)m2^{n}=\frac{1}{4}\cdot 8\cdot 256=512$. Hence
    $\mathbb{P}\bigl(\mathrm{Stab}_{\Gamma}(F)\neq\{\mathrm{id}\}\bigr)\;\leq\;2^{144-512}=2^{-368}.$
    This shows that a uniformly random $8$-bit S-box has a nontrivial EA-stabilizer with probability at most $2^{-368}$, an astronomically small quantity. The collision probability for EA-equivalence is similarly bounded by $2^{-mq^{n}+O(n^{2}+m^{2})}=2^{-2048+O(128)}\approx 2^{-1900}$.
Thus, for dimensions $n\geq 8$ (the regime relevant to modern block ciphers such as AES), the trivial-stabilizer property holds with overwhelming probability.
Remark 3.6.
Our upper bounds on $|\mathrm{Fix}(g)|$ and the resulting estimate $\mathbb{P}[\mathrm{Stab}(F)\neq\{\mathrm{id}\}]\leq q^{-\Omega(mq^{n})}$ are obtained via worst-case analysis. In principle, one could refine the argument by exploiting the cycle structure of affine permutations on $U$ and $W$ and by computing the rank distribution of QIQ-I for $Q\in\mathrm{GL}(W)$. Such refinements would slightly improve the constants in the exponent, but would not affect the asymptotic type of the decay, and would require a substantial amount of additional case analysis. For the purposes of establishing asymptotic freeness, the present bounds are entirely sufficient.
Corollary 3.7 (Asymptotic count of EA-classes).
Denote the set of EA-classes by $\mathcal{F}/\Gamma=\{\,\Gamma\cdot F:F\in\mathcal{F}\,\}$. Then
$\frac{|\mathcal{F}/\Gamma|}{|\mathcal{F}|/|\Gamma|}=1+o(1).$
In other words, the relative error between the true number of EA-classes and the naive estimate $|\mathcal{F}|/|\Gamma|$ is $o(1)$.
Proof.
Burnside’s lemma gives
$|\mathcal{F}/\Gamma|=\frac{1}{|\Gamma|}\sum_{g\in\Gamma}|\mathrm{Fix}(g)|.$
The identity contributes $|\mathcal{F}|$. Each nontrivial $g$ contributes at most $q^{cmq^{n}}$ by Lemma 3.3. Hence
$|\mathcal{F}/\Gamma|=\frac{|\mathcal{F}|}{|\Gamma|}+O\!\left(|\Gamma|\cdot q^{cmq^{n}}\right).$
The relative error is
$\frac{|\Gamma|\cdot q^{cmq^{n}}}{|\mathcal{F}|/|\Gamma|}=\frac{|\Gamma|^{2}\cdot q^{cmq^{n}}}{|\mathcal{F}|}=q^{-(1-c)mq^{n}+O(n^{2}+m^{2})}=o(1).$
To justify the exponent, note that $|\mathrm{AGL}(U)|=q^{n^{2}+O(n)}$ and $|\mathrm{AGL}(W)|=q^{m^{2}+O(m)}$, so $|\Gamma|=q^{n^{2}+m^{2}+O(n+m)}$, hence $|\Gamma|^{2}=q^{2n^{2}+2m^{2}+O(n+m)}=q^{O(n^{2}+m^{2})}$. Thus the relative error tends to zero. ∎
For the binary case ($q=2$), we have
$\frac{|\mathcal{F}/\Gamma|\cdot|\mathrm{AGL}(n,2)|\cdot|\mathrm{AGL}(m,2)|}{2^{m2^{n}}}=1+o(1).$
Consequently, the contribution of functions with nontrivial EA-stabilizers to the total count is exponentially negligible: their number is exponentially smaller than the full space. This observation is particularly relevant for structured searches in the design of cryptographic S-boxes.
Proposition 3.8 (Upper bounds on collision probabilities).
For two independently and uniformly random functions $F,G\in\mathcal{F}$, the probability that they are equivalent satisfies:
  1. 1.
    EA-equivalence:
    $\mathbb{P}(F\sim_{\mathrm{EA}}G)\;\leq\;\frac{|\mathrm{AGL}(U)|\cdot|\mathrm{AGL}(W)|}{|\mathcal{F}|}\;=\;q^{-mq^{n}+O(n^{2}+m^{2})}\;=\;q^{-\Omega(q^{n})}.$
  2. 2.
    CCZ-equivalence:
    $\mathbb{P}(F\sim_{\mathrm{CCZ}}G)\;\leq\;\frac{|\mathrm{AGL}(U\times W)|}{|\mathcal{F}|}\;=\;q^{-mq^{n}+O((n+m)^{2})}\;=\;q^{-\Omega(q^{n})}.$
Proof.
For both equivalence relations, the collision probability can be written as
$\mathbb{P}(F\sim G)=\sum_{i=1}^{r}\Bigl(\frac{|O_{i}|}{|\mathcal{F}|}\Bigr)^{2}\leq\sum_{i=1}^{r}\frac{|O_{i}|}{|\mathcal{F}|}\cdot\frac{\max_{j}|O_{j}|}{|\mathcal{F}|}=\frac{\max_{j}|O_{j}|}{|\mathcal{F}|},$
where $O_{1},\dots,O_{r}$ are the orbits. By the orbit-stabilizer theorem, the maximum orbit size is bounded by the group order. For EA-equivalence, $\max_{j}|O_{j}|\leq|\Gamma|=|\mathrm{AGL}(U)|\cdot|\mathrm{AGL}(W)|$, and for CCZ-equivalence, $\max_{j}|O_{j}|\leq|\mathrm{AGL}(U\times W)|$. The size estimates follow from
$|\mathrm{AGL}(n,q)|=q^{n}\cdot|\mathrm{GL}(n,q)|=q^{n^{2}+O(n)}$
and $|\mathcal{F}|=q^{mq^{n}}$, giving the stated bounds. ∎
The upper bounds in Proposition 3.8 follow directly from the orbit-stabilizer theorem and hold for both EA- and CCZ-equivalence over arbitrary finite fields. However, using our result on trivial stabilizers (Theorem 3.4), we can show that the EA upper bound is asymptotically tight:
Corollary 3.9 (Asymptotic collision probability for EA-equivalence).
For two independently and uniformly random functions $F,G\in\mathcal{F}$, the probability that $F$ and $G$ are EA-equivalent satisfies
$\mathbb{P}(F\sim_{\mathrm{EA}}G)=\frac{|\mathrm{AGL}(U)|\cdot|\mathrm{AGL}(W)|}{|\mathcal{F}|}\,(1+o(1)).$
Proof.
Let $O_{1},\dots,O_{r}$ be the EA-orbits. The collision probability is
$\mathbb{P}(F\sim_{\mathrm{EA}}G)=\sum_{i=1}^{r}\Bigl(\frac{|O_{i}|}{|\mathcal{F}|}\Bigr)^{2}=\frac{1}{|\mathcal{F}|^{2}}\sum_{i=1}^{r}|O_{i}|^{2}.$ (3)
Let $\mathcal{O}_{\mathrm{triv}}$ be the set of orbits with trivial EA-stabilizer and $\mathcal{O}_{\mathrm{nt}}$ the set of the remaining orbits. By Theorem 3.4, for every $\varepsilon>0$ and all sufficiently large $n$,
$|\{F\in\mathcal{F}:\mathrm{Stab}_{\Gamma}(F)=\{\mathrm{id}\}\}|\geq(1-\varepsilon)|\mathcal{F}|.$ (4)
Hence
$\sum_{O_{i}\in\mathcal{O}_{\mathrm{triv}}}|O_{i}|=|\{F:\mathrm{Stab}_{\Gamma}(F)=\{\mathrm{id}\}\}|\geq(1-\varepsilon)|\mathcal{F}|.$
Since each such orbit has size $|\Gamma|$, this implies
$|\mathcal{O}_{\mathrm{triv}}|\geq\frac{(1-\varepsilon)|\mathcal{F}|}{|\Gamma|}.$
The contribution of the trivial-stabilizer orbits to (3) is therefore
$\frac{1}{|\mathcal{F}|^{2}}\sum_{O_{i}\in\mathcal{O}_{\mathrm{triv}}}|O_{i}|^{2}=\frac{|\mathcal{O}_{\mathrm{triv}}|\cdot|\Gamma|^{2}}{|\mathcal{F}|^{2}}\geq(1-\varepsilon)\frac{|\Gamma|}{|\mathcal{F}|}.$
For the nontrivial-stabilizer orbits, since $|O_{i}|\leq|\Gamma|$ and $\sum_{O_{i}\in\mathcal{O}_{\mathrm{nt}}}|O_{i}|\leq\varepsilon|\mathcal{F}|$ by (4), we have
$\frac{1}{|\mathcal{F}|^{2}}\sum_{O_{i}\in\mathcal{O}_{\mathrm{nt}}}|O_{i}|^{2}\leq\varepsilon\frac{|\Gamma|}{|\mathcal{F}|}.$
Combining both contributions with the upper bound from Proposition 3.8 yields
$(1-\varepsilon)\frac{|\Gamma|}{|\mathcal{F}|}\;\leq\;\mathbb{P}(F\sim_{\mathrm{EA}}G)\;\leq\;\frac{|\Gamma|}{|\mathcal{F}|}.$
Since $\varepsilon>0$ is arbitrary, the result follows. ∎
This asymptotic equality validates random sampling strategies for searching cryptographic primitives. For two independently sampled functions, the probability of landing in the same EA-equivalence class is negligibly small. In the binary case, the collision probability is $2^{-\Omega(2^{n})}$, which means that random generation of vectorial Boolean functions produces EA-inequivalent candidates with overwhelming probability. This is particularly useful for exhaustive searches of functions with special cryptographic properties such as almost perfect nonlinear (APN) and almost bent (AB) functions: researchers can sample the function space randomly without concern for redundancy modulo EA-equivalence, as collisions are exponentially rare.
4 Conclusion
We have established that asymptotically almost all vectorial functions over finite fields have trivial EA-stabilizers. This fundamental result resolves the two questions posed in the introduction.
  1. 1.
    Classification: The number of EA-equivalence classes is asymptotically equal to the naive estimate (the total number of functions divided by the size of the EA-group) with vanishing relative error. This shows that the EA group action is asymptotically free: almost every function has a trivial EA-stabilizer. For cryptographic applications, this means that the vast majority of vectorial functions lie in distinct EA-classes, and functions with nontrivial EA-stabilizers form an exponentially small subset.
  2. 2.
    Random sampling: Together with upper bounds on collision probabilities, our results establish that for EA-equivalence, two independently sampled functions are equivalent with super-exponentially small probability. This validates random sampling strategies for searching cryptographic primitives such as APN or AB functions: exhaustive random searches produce EA-inequivalent candidates with overwhelming probability, eliminating concerns about redundancy modulo EA-equivalence.
While we have proved upper bounds on collision probabilities for CCZ-equivalence, the question of whether most functions have trivial CCZ-stabilizers remains open. Establishing analogous results for CCZ-equivalence would require new techniques, as the graph-based action is more complex than the EA group action.